OpenStudy (anonymous):

What number decreased by 1/5 if itself equals 84?

OpenStudy (anonymous):

x-(1/5)x=84 (4/5)x=84 x=(5/4)*84=21*5=105

OpenStudy (anonymous):

Thanks